Please help me with this question:
The length of a rectangular playing field is 5m longer than its width. If the perimeter of the field is 150m, find its width.

Answer by Alan3354(69443)

Please help me with this question:
The length of a rectangular playing field is 5m longer than its width. If the perimeter of the field is 150m, find its width.
--------------
Perimeter = 2L + 2W = 150
L + W = 75
L - W = 5
------------------ Add